# Negative Dependence as a toolbox for machine learning : review and new developments

### v1

@@ Abstract
[…]
@@ 1.3 Related literature
Random point sets with negative dependence have been in vogue as a scientific discipline for several decades. The principal motivation for this came from statistical and quantum physics. An early connection arose in the context of Fermionic particle systems, wherein determinantal structures were observed in ground state densities of Fermionic systems via Slater determinants [Sla29]. Subsequently, it was developed as a mathematical discipline by Macchi and others [Mac72, Mac75, Sos00], and new connections arose with a wide array of areas in mathematics and statistical physics, including random matrices and Coulomb gases [Joh05, Bor15], integrable systems [Dei08, Dei17], combinatorics [BG15, BP16], among others.
@@
Negative dependence as a tool for machine learning and statistics is of a more recent historical provenance. Earlier work in this direction is well explored in the excellent survey by Kulesza and Taskar [KT12], which appeared only around a decade ago, where we refer the interested reader for an account of formative developments. This was also roughly contemporaneous with the emergence of works in the statistical literature that looked at DPPs and allied negatively dependent processes as a statistical modelling paradigm for strongly correlated data [LMR14, LPW21, PL23]. These developments motivated investigations into statistical estimation and inference on DPPs, e.g., learning the parameters of DPPs [AFAT14, UBMR17], efficient sampling schemes for continuous DPPs [AFT13], rates of estimation for DPPs [BMRU17], and DPPs for minibatch sampling in stochastic gradient descent [BGL21].
@@
This also ties in with a burgeoning interest, principally in the theoretical computer science and adjacent communities, in carrying out investigations into general models of negatively associated random sets (in the discrete setting), known as strongly Rayleigh measures. Within the limited boundaries of the present article, we will not have the occasion to delve into this beautiful theory, instead referring the reader to the seminal works of Borcea, Branden and Liggett [BB09a, BB09b, BB09c, BBL09, Br 07, Br 14], for the development of the general theory as well as its connections to other parts of discrete mathematics, such as matroid theory, log concavity of sequences and related topics. In the setting of applications to learning theory, notable contributions include counting bases of matroids [ALGV24], sampling algorithms [AOGR16, LJS16], traveling salesman problem [AGM+17], learning with DPPs [GHD+20, MS15, MS17]. On a related note, we must mention the resolution of the Kadison-Singer conjecture and construction of a very general class of Ramanujan graphs by Marcus, Spielman and Srivastava over the last decade that strongly leverages zeros of characteristic polynomials and their interlacing structures [MSS15a, MSS15b, MSS22, MSS18].
@@ 2.1 Basic notions
Determinantal point processes (DPPs) are random configurations of locally finite point sets over some background space, whose correlation functions are given by determinants of certain kernels. They have a long history in both mathematics and physics. DPPs were first introduced by Macchi [Mac72], motivated by fermions in quantum mechanics. Macchi discovered that DPPs describe the distribution of a fermionic system at thermal equilibrium; their repulsive behavior captures the Pauli exclusion principle, which states that two fermions cannot occupy the same quantum state. Surprisingly, DPPs also arise naturally in many mathematical settings, such as the eigenvalues of random matrices, random spanning trees, and zeros of random analytic functions, among others. As such, DPPs have intertwined with various fields in mathematics, including random matrix theory [Bor15, Joh05], integrable systems [Dei08, Dei17], combinatorics [BG15, BP16], and complex geometry [Ber20, Ber13, Ber14, Ber21], to name a few.
@@
Under mild conditions, the joint distribution of $\Lambda_{\mathcal{S}}(f)$ for a sufficiently rich class of test functions $f$ (e.g., the class of all continuous, compactly supported functions) can determine the distribution of the process $\mathcal{S}$; thus, studying linear statistics is of fundamental interest. For DPPs, due to the determinantal structure, there is an explicit formula for the Laplace transform of linear statistics. As a consequence, one can deduce explicit formulas for the cumulants (hence, moments) of linear statistics of DPPs (e.g., see [JL18]).
@@
We warn the readers that not every kernel $K$ defines a DPP. In fact, we have the following theorem by Macchi [Mac75] and Soshnikov [Sos00].
@@
An important subclass of DPPs, called projection DPPs, consists of those for which the associated integral operator $\mathcal{K}:L^{2}(\mathcal{X},\mu)\rightarrow L^{2}(\mathcal{X},\mu)$ is self-adjoint, bounded, and has all non-zero eigenvalues equal to 1 (equivalently, $\mathcal{K}$ is a projection operator onto a subspace of $L^{2}(\mathcal{X},\mu)$). It turns out that all DPPs with Hermitian kernels are mixtures of projection DPPs (see [HKPV05]).
@@
Let $\mathcal{S}$ be a DPP on $\mathbb{R}^{d}$ with Hermitian, translation-invariant kernel $K$ (i.e., $K(x,y)=\overline{K(y,x)}$ and $K(x,y)=\Phi(x-y)$ for some $\Phi$), and background measure the Lebesgue measure. Let $B_{\varepsilon}$ be a ball of small radius $\varepsilon>0$ (one can assume the center is 00 due to the translation invariance) and denote by $\mathcal{S}(B_{\varepsilon})$ the number of points in $\mathcal{S}$ inside $B_{\varepsilon}$. Then we have
@@
For more detailed discussion and further properties of DPPs, we refer the readers to [Sos00, HKPV05, Lyo03, Lyo14] for excellent overviews.
@@ 2.2 DPPs in machine learning
In recent years, DPPs have attracted the attention of the machine learning community. They are generic statistical models for repulsion in spatial statistics [LMR14, BL17] and machine learning [KT12, BBC20, GBDK19, GHD+20, DM19, TBA19, GR20]. Sampling and inference with DPPs are tractable, and there are efficient algorithms for such tasks [KT12, GR20, HGG+22, HGDK22].
@@
where $\mathbf{L}:=\mathbf{K}(\mathbf{I}-\mathbf{K})^{-1}$ and $\mathbf{L}_{A}$ denotes the submatrix of $\mathbf{L}$ whose rows and columns indexed by the subset $A$ (see [KT12]).
In general, the cardinality $|\mathcal{S}|$ of a DPP $\mathcal{S}$ is an $\mathbb{N}$-valued random variable. In applications, sometimes it is interesting to sample a point process similarly to a DPP but with fixed size $m$ (called an $m$-DPP in [KT12]). This can be done by a conditioning argument as follows. Given a DPP $\mathcal{S}$ on $\mathcal{X}$, one can obtain an $m$-DPP by considering the conditioned process $(\mathcal{S}|\{|\mathcal{S}|=m\})$, provided that $\mathbb{P}(|\mathcal{S}|=m)>0$. However, $m$-DPPs do not have the determinantal structure in general (i.e., they are not DPPs); the only exception is when the original DPP $\mathcal{S}$ is a projection DPP.
@@ 3 Sampling from DPPs
DPPs are particularly useful tools for machine learning since they can be sampled in polynomial time. The classic algorithm to sample a DPP comes from [HKPV05] which leverages a neat geometric understanding of a DPP as a probability measure. Recent algorithmic advances significantly improve this classical algorithm to achieve computational efficiency.
@@ 3.1 The classic Spectral Sampler for DPPs
[…]
@@
[…]
@@
This algorithm is an avatar of the complete algorithm of [HKPV05], and the proof of validity in the discrete form we use here can be found in [KT12]. When $N$ becomes large, the true bottleneck here is accessing the eigendecomposition of $\mathbf{L}$, which required $\mathcal{O}(N^{3})$ operations. For kernels with low rank $r\ll N$, the computational cost of the eigendecomposition reduces to $\mathcal{O}(r^{2}N)$ by using dual kernel.
@@ 3.2 Tree-based algorithms for computational efficiency
When repeated DPP samples are required, economy can be achieved by reusing the eigendecomposition of the kernel, to be computed in the preprocessing step. For the second phase, given the eigendecomposition, the complexity of getting each additional sample is linear in $N$ in Algorithm 2. This dependence can be significantly reduced to $\log N$ by using a tree-based algorithm for fast repeated sampling from DPPs, developed recently in [GKMV19].
To elaborate, let $\mathcal{S}$ be an $L$-ensemble with kernel $\mathbf{L}\in\mathbb{R}^{N\times N}$ of rank $r$. We write $\mathbf{L}=\mathbf{B}^{\top}\mathbf{B}$, where $\mathbf{B}\in\mathbb{R}^{r\times N}$ and define $\mathbf{C}:=\mathbf{B}\mathbf{B}^{\top}\in\mathbb{R}^{r\times r}$, called the dual kernel. In practice, when the user provides $\mathbf{B}$, the dual kernel $\mathbf{C}$ is preferred in sampling low-rank DPPs, as it reduces the eigendecomposition cost to $\mathcal{O}(r^{2}N)$. We consider the eigendecomposition of $\mathbf{C}=\sum_{i=1}^{r}\lambda_{i}\mathbf{w}_{i}\mathbf{w}_{i}^{\top}$. The first phase of the tree-based algorithm in [GKMV19] is the same as the first phase in Algorithm 2, where we use the eigendecomposition to sample a subset of indices $E\subset\{1,2,\ldots,r\}$ by independently picking $i\in E$ with probability $\lambda_{i}/(\lambda_{i}+1)$. The advancement comes in the second phase, where instead of directly sampling from a multinomial distribution defined by a projection kernel (obtained in the first phase) and repeatedly updating the kernel as in Algorithm 2 (which is costly), we will repeatedly sample from a binary tree of depth $\log N$. The tree construction, appropriate summary statistics stored in the tree will be done in the pre-processing phase, which is a one-time cost.
@@
6:$S_{l},S_{r}\leftarrow\mathrm{Split}(S)$
7:$\mathsf{T}_{l}\leftarrow\mathrm{ConstructTree}(S_{l},\gamma,\mathbf{G},\mathbf{H})$
8:$\mathsf{T}_{r}\leftarrow\mathrm{ConstructTree}(S_{r},\gamma,\mathbf{G},\mathbf{H})$
9:$\mathbf{z}\leftarrow\mathsf{T}_{l}.\mathbf{z}+\mathsf{T}_{r}.\mathbf{z}$
10:$\mathbf{A}\leftarrow\mathsf{T}_{l}.\mathbf{A}+\mathsf{T}_{r}.\mathbf{A}$
@@
Proposition 3.2 (Proposition 1 in [GKMV19]).
@@
13:$p_{l}\leftarrow\mathrm{ComputeMarginal}(\mathsf{T}_{l},E,Y,\mathbf{G},\mathbf{Q})$
14:$p_{r}\leftarrow\mathrm{ComputeMarginal}(\mathsf{T}_{r},E,Y,\mathbf{G},\mathbf{Q})$
@@
More recently, many sophisticated sampling algorithms for DPPs have been explored and developed, extending the tree-based approach or otherwise. However, within the limited scope of this article we only address the most fundamental techniques, referring the interested reader to [HGG+22, HGDK22, GKMV19, BTA22, ACV24] and the references therein for recent developments.
@@ 4.2 Existing algorithms
Given the classical nature of the problem, many quadrature methods have been developed, including Gaussian quadrature [DR84, Gau04, BP11, GV83], Monte Carlo methods [RC05, BDH17], Quasi-Monte Carlo methods [DP10, DKS13, OG16], Monte Carlo with DPPs [BH18, BH20, GBV19, LB24] and Bayesian quadrature [HD12, BOGO15], to name a few. In this survey, we briefly review Gaussian quadrature, Monte Carlo methods and Monte Carlo with DPPs.
@@ 4.2.1 Gaussian quadrature
We consider the case when $\mu$ is a measure supported on the interval $I:=[-1,1]$. Let $(\varphi_{k})_{k\in\mathbb{N}}$ be a sequence of polynomials such that for each $k\in\mathbb{N}$, $\varphi_{k}$ has degree $k$, positive leading coefficient, and such that
@@
Gaussian quadrature [DR84, Gau04, BP11] outputs $N$ nodes $x_{1},\ldots,x_{N}$ to be the zeros of $\varphi_{N}(x)$ (which are real and simple) and weights $w_{i}=K_{N}(x_{i},x_{i})^{-1}$. The error is then given by
@@
Gaussian quadrature is particularly effective when the test function $f$ is close to a polynomial. Specifically, this method gives $\mathcal{E}_{N}(f)=0$ for every polynomial $f$ of degree less than $12N-$ 1, and $\mathcal{E}_{N}(f)$ decays exponentially fast when $f$ is analytic (e.g., see [GV83]). However, optimal rates of decay for $\mathcal{E}_{N}(f)$ are still unknown when $f$ is less regular. For instance, when $f$ is $C^{1}$, it is only known that $\mathcal{E}_{N}(f)=O(N^{-1})$, based on the classical Jackson’s approximation theorem.

### v4

@@ Abstract
$\dagger$$\dagger$}\$footnotetext: Corresponding author\$1\$\$1\$footnotetext: Department of Mathematics, National University of Singapore, 10 Lower Kent Ridge Road, 119076, Singapore\$2\$\$2\$footnotetext: Univ. Lille, CNRS, Centrale Lille, UMR 9189 – CRIStAL, F-59000 Lille, France
@@ 1.3 Related literature
Random point sets with negative dependence have been in vogue as a scientific discipline for several decades. The principal motivation for this came from statistical and quantum physics. An early connection arose in the context of Fermionic particle systems, wherein determinantal structures were observed in ground state densities of Fermionic systems via Slater determinants [Slater]. Subsequently, it was developed as a mathematical discipline by Macchi and others [Mac72, Mac75, SoshDPP], and new connections arose with a wide array of areas in mathematics and statistical physics, including random matrices and Coulomb gases [johanssondpp, borodin], integrable systems [Deift_1, Deift_2], combinatorics [borodin2015_integrable, borodin2016_integrable], among others.
@@
Negative dependence as a tool for machine learning and statistics is of a more recent historical provenance. Earlier work in this direction is well explored in the excellent survey by Kulesza and Taskar [kulesza_determinantal_2012], which appeared only around a decade ago, where we refer the interested reader for an account of formative developments. This was also roughly contemporaneous with the emergence of works in the statistical literature that looked at DPPs and allied negatively dependent processes as a statistical modelling paradigm for strongly correlated data [LaMoRu14, lavancier-1, lavancier-2]. These developments motivated investigations into statistical estimation and inference on DPPs, e.g., learning the parameters of DPPs [Fox-Affandi-1, Rigollet-2], efficient sampling schemes for continuous DPPs [Fox-Affandi-2], rates of estimation for DPPs [Rigollet-1], and DPPs for minibatch sampling in stochastic gradient descent [OPE-NIPS].
@@
This also ties in with a burgeoning interest, principally in the theoretical computer science and adjacent communities, in carrying out investigations into general models of negatively associated random sets (in the discrete setting), known as strongly Rayleigh measures. Within the limited boundaries of the present article, we will not have the occasion to delve into this beautiful theory, instead referring the reader to the seminal works of Borcea, Branden and Liggett [Borcea-Branden, Borcea-Branden-2, Borcea-Branden-3, Borcea-Branden-Liggett, BRANDEN, Branden-2], for the development of the general theory as well as its connections to other parts of discrete mathematics, such as matroid theory, log concavity of sequences and related topics. In the setting of applications to learning theory, notable contributions include counting bases of matroids [anari_logconcave2], sampling algorithms [anari_mcmc_rayleigh_dpp, Jegelka], traveling salesman problem [travelling-salesman], learning with DPPs [gartrell2020scalable, Sra, mariet_diversity_2017]. On a related note, we must mention the resolution of the Kadison-Singer conjecture and construction of a very general class of Ramanujan graphs by Marcus, Spielman and Srivastava over the last decade that strongly leverages zeros of characteristic polynomials and their interlacing structures [interlacing-1, Marcus-Spielman-Srivastava, interlacing-3, interlacing-4].
@@ 2.1 Basic notions
Determinantal point processes (DPPs) are random configurations of locally finite point sets over some background space, whose correlation functions are given by determinants of certain kernels. They have a long history in both mathematics and physics. DPPs were first introduced by Macchi [Mac72], motivated by fermions in quantum mechanics. Macchi discovered that DPPs describe the distribution of a fermionic system at thermal equilibrium; their repulsive behavior captures the Pauli exclusion principle, which states that two fermions cannot occupy the same quantum state. Surprisingly, DPPs also arise naturally in many mathematical settings, such as the eigenvalues of random matrices, random spanning trees, and zeros of random analytic functions, among others. As such, DPPs have intertwined with various fields in mathematics, including random matrix theory [borodin, johanssondpp], integrable systems [Deift_1, Deift_2], combinatorics [borodin2015_integrable, borodin2016_integrable], and complex geometry [Berman0, Berman1, Berman2, Berman3], to name a few.
@@
Under mild conditions, the joint distribution of $\Lambda_{\mathcal{S}}(f)$ for a sufficiently rich class of test functions $f$ (e.g., the class of all continuous, compactly supported functions) can determine the distribution of the process $\mathcal{S}$; thus, studying linear statistics is of fundamental interest. For DPPs, due to the determinantal structure, there is an explicit formula for the Laplace transform of linear statistics. As a consequence, one can deduce explicit formulas for the cumulants (hence, moments) of linear statistics of DPPs (e.g., see [lambertdpps]).
@@
We warn the readers that not every kernel $K$ defines a DPP. In fact, we have the following theorem by Macchi [Mac75] and Soshnikov [SoshDPP].
@@
An important subclass of DPPs, called projection DPPs, consists of those for which the associated integral operator $\mathcal{K}:L^{2}(\mathcal{X},\mu)\rightarrow L^{2}(\mathcal{X},\mu)$ is self-adjoint, bounded, and has all non-zero eigenvalues equal to 1 (equivalently, $\mathcal{K}$ is a projection operator onto a subspace of $L^{2}(\mathcal{X},\mu)$). It turns out that all DPPs with Hermitian kernels are mixtures of projection DPPs (see [HKPV]).
@@
Let $\mathcal{S}$ be a DPP on $\mathbb{R}^{d}$ with Hermitian, translation-invariant kernel $K$ (i.e., $K(x,y)=\overline{K(y,x)}$ and $K(x,y)=\Phi(x-y)$ for some $\Phi$), and background measure the Lebesgue measure. Let $B_{\varepsilon}$ be a ball of small radius $\varepsilon>0$ (one can assume the center is 0 due to the translation invariance) and denote by $\mathcal{S}(B_{\varepsilon})$ the number of points in $\mathcal{S}$ inside $B_{\varepsilon}$. Then we have
@@
For more detailed discussion and further properties of DPPs, we refer the readers to [SoshDPP, HKPV, LyonsIHES, lyons2014survey] for excellent overviews.
@@ 2.2 DPPs in machine learning
In recent years, DPPs have attracted the attention of the machine learning community. They are generic statistical models for repulsion in spatial statistics [LaMoRu14, biscio2017contrast] and machine learning [kulesza_determinantal_2012, belhadji_determinantal_2020, gartrell2019learning, gartrell2020scalable, determinantal-averaging, tremblay2019determinantal, GDP]. Sampling and inference with DPPs are tractable, and there are efficient algorithms for such tasks [kulesza_determinantal_2012, GDP, gartrell_scalablesampling, han2022scalable].
@@
where $\mathbf{L}:=\mathbf{K}(\mathbf{I}-\mathbf{K})^{-1}$ and $\mathbf{L}_{A}$ denotes the submatrix of $\mathbf{L}$ whose rows and columns indexed by the subset $A$ (see [kulesza_determinantal_2012]).
In general, the cardinality $|\mathcal{S}|$ of a DPP $\mathcal{S}$ is an $\mathbb{N}$-valued random variable. In applications, sometimes it is interesting to sample a point process similarly to a DPP but with fixed size $m$ (called an $m$-DPP in [kulesza_determinantal_2012]). This can be done by a conditioning argument as follows. Given a DPP $\mathcal{S}$ on $\mathcal{X}$, one can obtain an $m$-DPP by considering the conditioned process $(\mathcal{S}|\{|\mathcal{S}|=m\})$, provided that $\mathbb{P}(|\mathcal{S}|=m)>0$. However, $m$-DPPs do not have the determinantal structure in general (i.e., they are not DPPs); the only exception is when the original DPP $\mathcal{S}$ is a projection DPP.
@@ 3 Sampling from DPPs
DPPs are particularly useful tools for machine learning since they can be sampled in polynomial time. The classic algorithm to sample a DPP comes from [HKPV] which leverages a neat geometric understanding of a DPP as a probability measure. Recent algorithmic advances significantly improve this classical algorithm to achieve computational efficiency.
@@ 3.1 The classic Spectral Sampler for DPPs
Let $\mathcal{S}$ be a projection DPP of rank $m$, i.e., its kernel $K$ defines a projection operator $\mathcal{K}_{H}$ from $L^{2}(\mathcal{X},\mu)$ to an $m$-dimensional subspace $H\subset L^{2}(\mathcal{X},\mu)$. We denote by $\|\cdot\|$ the norm in the Hilbert space $L^{2}(\mathcal{X},\mu)$. Then for each $f\in L^{2}(\mathcal{X},\mu)$, $\mathcal{K}_{H}f$ is the unique element in $H$ closest to $f$.
@@
Let $(X_{1},\ldots,X_{m})$ be the points constructed by Algorithm 1. Then $\mathcal{S}\overset{d}{=}\{X_{1},\ldots,X_{m}\}$.
@@
This algorithm is an avatar of the complete algorithm of [HKPV], and the proof of validity in the discrete form we use here can be found in [kulesza_determinantal_2012]. When $N$ becomes large, the true bottleneck here is accessing the eigendecomposition of $\mathbf{L}$, which required $\mathcal{O}(N^{3})$ operations. For kernels with low rank $r\ll N$, the computational cost of the eigendecomposition reduces to $\mathcal{O}(r^{2}N)$ by using dual kernel.
@@ 3.2 Tree-based algorithms for computational efficiency
When repeated DPP samples are required, economy can be achieved by reusing the eigendecomposition of the kernel, to be computed in the preprocessing step. For the second phase, given the eigendecomposition, the complexity of getting each additional sample is linear in $N$ in Algorithm 2. This dependence can be significantly reduced to $\log N$ by using a tree-based algorithm for fast repeated sampling from DPPs, developed recently in [kulesza-gillenwater].
To elaborate, let $\mathcal{S}$ be an $L$-ensemble with kernel $\mathbf{L}\in\mathbb{R}^{N\times N}$ of rank $r$. We write $\mathbf{L}=\mathbf{B}^{\top}\mathbf{B}$, where $\mathbf{B}\in\mathbb{R}^{r\times N}$ and define $\mathbf{C}:=\mathbf{B}\mathbf{B}^{\top}\in\mathbb{R}^{r\times r}$, called the dual kernel. In practice, when the user provides $\mathbf{B}$, the dual kernel $\mathbf{C}$ is preferred in sampling low-rank DPPs, as it reduces the eigendecomposition cost to $\mathcal{O}(r^{2}N)$. We consider the eigendecomposition of $\mathbf{C}=\sum_{i=1}^{r}\lambda_{i}\mathbf{w}_{i}\mathbf{w}_{i}^{\top}$. The first phase of the tree-based algorithm in [kulesza-gillenwater] is the same as the first phase in Algorithm 2, where we use the eigendecomposition to sample a subset of indices $E\subset\{1,2,\ldots,r\}$ by independently picking $i\in E$ with probability $\lambda_{i}/(\lambda_{i}+1)$. The advancement comes in the second phase, where instead of directly sampling from a multinomial distribution defined by a projection kernel (obtained in the first phase) and repeatedly updating the kernel as in Algorithm 2 (which is costly), we will repeatedly sample from a binary tree of depth $\log N$. The tree construction, appropriate summary statistics stored in the tree will be done in the pre-processing phase, which is a one-time cost.
@@
[…]
@@
Proposition 3.2 (Proposition 1 in [kulesza-gillenwater]).
@@
[…]
@@
More recently, many sophisticated sampling algorithms for DPPs have been explored and developed, extending the tree-based approach or otherwise. However, within the limited scope of this article we only address the most fundamental techniques, referring the interested reader to [gartrell_scalablesampling, han2022scalable, kulesza-gillenwater, Tremblay-samplingdpp, anari-samplingdpp] and the references therein for recent developments.
@@ 4.2 Existing algorithms
Given the classical nature of the problem, many quadrature methods have been developed, including Gaussian quadrature [DavisRabi, Gautschi, BrassPetras, GautschiVarga], Monte Carlo methods [MonteCarlobook, bardenet:MCMC], Quasi-Monte Carlo methods [Dick_Pillichshammer_2010, Dick_2013, Oates16], Monte Carlo with DPPs [bardenet:montecarlogaussian, OPE-AOAP, bardenet:montecarlodpp, lemoine2024montecarlomethodscompact] and Bayesian quadrature [Huszar12, Briol2015], to name a few. In this survey, we briefly review Gaussian quadrature, Monte Carlo methods and Monte Carlo with DPPs.
@@ 4.2.1 Gaussian quadrature
We consider the case when $\mu$ is a measure supported on the interval I:=[1,1]I:=[-1,1]. Let $(\varphi_{k})_{k\in\mathbb{N}}$ be a sequence of polynomials such that for each $k\in\mathbb{N}$, $\varphi_{k}$ has degree $k$, positive leading coefficient, and such that
@@
Gaussian quadrature [DavisRabi, Gautschi, BrassPetras] outputs $N$ nodes $x_{1},\ldots,x_{N}$ to be the zeros of $\varphi_{N}(x)$ (which are real and simple) and weights $w_{i}=K_{N}(x_{i},x_{i})^{-1}$. The error is then given by
@@
Gaussian quadrature is particularly effective when the test function $f$ is close to a polynomial. Specifically, this method gives $\mathcal{E}_{N}(f)=0$ for every polynomial $f$ of degree less than 2N12N-1, and $\mathcal{E}_{N}(f)$ decays exponentially fast when $f$ is analytic (e.g., see [GautschiVarga]). However, optimal rates of decay for $\mathcal{E}_{N}(f)$ are still unknown when $f$ is less regular. For instance, when $f$ is $C^{1}$, it is only known that $\mathcal{E}_{N}(f)=O(N^{-1})$, based on the classical Jackson’s approximation theorem.